OpenStudy (anonymous):

The sum of two numbers is 110. Their difference is 28. What is the smallest of the two numbers? A. 56 B. 41 C. 69 D. 55

OpenStudy (lgbasallote):

let x = bigger y = smaller x + y =110 x - y = 28 elimination... 2x = 138 x = 69 y = 110 - x y = 110 - 69 y = 41 smaller number is 41